Convert the numeral to a numeral in the base indicated.
16,629 to base 13

Answer by Alan3354(69443)

16,629 to base 13
-----------
13^3 = 2197
16629/2197 = 7 + remainder
rem = 1250
1250/169 = 7 + rem
67
67 = 5*13 + 2
--> 7752